40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 13 03:22:53 2011, kiwamu, Update, LSC, measurement of sensing matrix : just began 
    Reply  Fri Oct 14 04:35:45 2011, kiwamu, Update, LSC, locking tonight Screen_shot_2011-10-14_at_4.24.01.png
Message ID: 5665     Entry time: Fri Oct 14 04:35:45 2011     In reply to: 5659
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: locking tonight 

The lock of DRMI wasn't stable enough to measure the sensing matrix. Failed.

PRMI and SRMI were okay and in fact they could stay locked robustly for a long time.

I added a new option in the C1IFO_CONFIGURE screen so that one can choose Signal-Recycled Michelson in carrier resonant condition.

Screen_shot_2011-10-14_at_4.24.01.png

 Additionally the orthogonalization of the I-Q signals on REFL55 should be done because it hasn't been done.

 

ELOG V3.1.3-